CAL (General)
chihiro.kozakai - 19:10, Saturday 03 November 2018 (6843)
X-end calibration line injection test

Calibration line injection over night

The calibration line injection of 15 Hz was running over night and it was still running well in this morning. Attached figure is spectrum at 18/11/3 3:00:00.

TxPD1
noise floor -80 dBV
15 Hz +18 dBv
30 Hz -55 dV
45 Hz -45 dBV
60 Hz -65 dBV
Bump @ 4500 Hz. 5 dB against noise. Observed in Tx1 and ToAOM2. Cross talk is suspected ?


TxPD2
noise floor -88 dBV
15 Hz +18 dBV
30 Hz -61 dBV
45 Hz -55 dBV
60 Hz -77 dBV
The noise floor has slope in low frequency.

The working point @ start of the injection (11/2 midnight)

Rx 1.862 V
LPD 7.64 V
path1
gain 1.8 V (15.494 dB)
offset 5.6 V (To AOM 0.304 V)
Tx1 7.099 V
OFS1 4.976 V
path2
gain 1.8 V (15.494 dB)
offset 6.5 V (ToAOM 0.359 V)
Tx2 6.009 V
OFS2 6.336 V

The working point @ ~10 hrs later (11/3 morning)

LPD 7.72 V
ToAOM1 0.3 V
Tx1 8.4 V
OFS1 5.9 V
ToAOM2 0.3 V
Tx2 6.4 V
OFS2 7.0 V
Rx 2.1 V

So, we observed the drift of the working point.

Mine work

In the afternoon, we went to X-end. We did following:

1. Cable labeling

2. GCal installation planned place check

3. WSK withdrawal

4. The 18 V DC source is found to have lower value. To be fixed.

5. Cleanup

In addition, accelerometer and microphone were installed by Yokozawa-san. Be careful when talking around the X-end PCal.

Calibration line injection failed accidentally

In the evening after mine work, we found that the PCal was not working. Control was lost. The cause is to be investigated.

Calibration line injection restart

We restarted the 15 Hz calibration line injection around 11/3 5pm. We will keep injecting until Monday morning.

 

Chihiro, Yuki, Takahiro S.

Images attached to this report
6843_20181103105422_51.png